OpenStudy (anonymous):

Two circles have radii in a ratio of 1:2. What is the ratio of their areas? 1:2 1:4 1:8 1:2π 1:4π

OpenStudy (ghazi):

ratio of area= ratio of radius square= (1/2)^2=1/4

OpenStudy (anonymous):

thnx!

OpenStudy (ghazi):

;)